Work problem:
Normally, it takes Jeffery 1 hour 45 minutes (105 min or 7/4 hour) to mow and trim an average lawn. When Lupe worked with him, an average lawn only took an hour. How long would it take Lupe to mow and trim an average yard on her own?

Answer by stanbon(75887)

Normally, it takes Jeffery 1 hour 45 minutes (105 min or 7/4 hour) to mow and trim an average lawn. When Lupe worked with him, an average lawn only took an hour. How long would it take Lupe to mow and trim an average yard on her own?
-----
Jeff time = 7/4 hr/job ; rate = (4/7) job/hr
------------------
Together time = 1 hr/job ; rate = 1 job/hr
---------------------
Lupe time = x hr/job ; rate = 1/x job/hr
-----------------
Equation:
rate + rate = together rate
4/7 + 1/x = 1
----
4x + 7 = 7x
3x = 7
x = 7/3 hrs (Lepe's time alone)
===================================
Cheers,
Stan H.
====================